How do I go about solving -7=x/5
Answer by nyc_function(2741)

-7 = x/5
-7 * 5 = (x/5) * 5
-35 = x
THE PROOF:
Let x = -35 in the original question given.
-7 = -35/5
-7 = -7...It checks!!!
The answer is x = -35.